# Engine-fed Kilonovae (Mergernovae) - II. Radiation

### v1

@@ 1 Introduction
Kilonova emission has garnered significant attention in recent years, particularly since the detection of AT2017gfo in association with the binary-neutron-star (BNS) merger GW170817 (e.g. Arcavi et al., 2017; Chornock et al., 2017; Cowperthwaite et al., 2017; Drout et al., 2017; Evans et al., 2017; Gao et al., 2017a; Kasen et al., 2017; Kilpatrick et al., 2017; Nicholl et al., 2017; Shappee et al., 2017; Smartt et al., 2017; Tanvir et al., 2017; Villar et al., 2017; Metzger, 2019, and the references there in, for a review). Other kilonova candidates have been observed in association with short-duration gamma-ray bursts (GRBs), which are believed to be also of a neutron star merger origin (e.g. Tanvir et al., 2013; Berger et al., 2013; Fan et al., 2013b; Yang et al., 2015; Jin et al., 2016; Gao et al., 2015, 2017b). Most recently, some unusual GRB events, specifically GRB 211211A and GRB 230307A, have long duration but with kilonova-like emission signal associated (Rastinejad et al., 2022; Yang et al., 2022; Troja et al., 2022; Gao et al., 2022; Zhu et al., 2022; Yang et al., 2023; Gillanders et al., 2023; Levan et al., 2023). The James Webb Space Telescope’s mid-infrared follow-up of GRB 230307A revealed an emission line at 2.15 microns in the spectrum, which is interpreted as tellurium (Levan et al., 2023; Gillanders et al., 2023). Additionally, the rapidly decaying bolometric luminosity of the kilonova at the later stage supports the presence of lanthanide in the ejecta. This kilonova signal is hence generally believed to be powered by r-process radioactive heating (Yang et al., 2023).
Notably, the presence of a plateau in the X-ray afterglow light curve of GRB 230307A, detected by the Lobster Eye Imager for Astronomy (LEIA), a pathfinder of the Einstein Probe mission, indicates that the merger remnant is a magnetar (Sun et al., 2023; Ling et al., 2023). Previous studies have speculated that a post-merger magnetar could drive an engine-fed kilonova (mergernova), which is anticipated to be significantly brighter than a typical r-process kilonova (Yu et al., 2013; Metzger & Piro, 2014; Gao et al., 2015; Sarin et al., 2022; Wang et al., 2024a). Therefore, the detection of the engine-fed kilonova signal serves as another crucial piece of evidence regarding the production of a long-lived magnetar. In reality, for some GRB events, the associated kilonova emission can be explained under both the r-process-driven and the magnetar-driven scenarios. For instance, AT2019gfo can be successfully fitted using the engine-fed kilonova (mergernova) model (Li et al., 2018; Yu et al., 2018), despite many believing it to be solely powered by r-process nucleosynthesis. This is because the large free parameter space can, to some extent, mitigate the differences between the engine-fed kilonova and the r-process one. Therefore, precise models are imperative for effectively distinguishing between them.
@@
We have made continuous efforts to more precisely model the engine-fed kilonova. Our first paper in this series (Ai et al., 2022) provided a detailed study for the dynamics of the ejecta resulting from the interaction between the magnetar wind and the merger ejecta. It is under the framework of the mechanical model for magnetized relativistic blastwaves (Ai & Zhang, 2021). For convenience, we henceforth refer to Ai et al. (2022) as Paper I. In Paper I, we found that a significant fraction ($\xi$ as high as $\sim 0.6$) of the energy from the post-merger magnetar’s wind can be converted into the kinetic energy of the ejecta’s bulk motion. Meanwhile, the forward shock excited in the ejecta provides an additional heating source during its propagation, with a heating efficiency of $\xi_{t}\sim(0.006-0.3)$. In this work, building upon the dynamical model proposed in Paper I, we focus on radiation. The basic formula for modeling both the heating and radiation processes of the engine-fed kilonova is presented in Section 2. The results, including the multi-band light curves and the spectral energy distribution (SED), are shown in Section 3. The conclusion and discussion are detailed in Section 4.
@@ 2 Methods
In this paper, focusing on the engine-fed kilonovae, we only use some simple analytical formulae to approximately describe the radioactive heating. Also, to simplify this problem, we simulate the heat conduction process in the radiation-dominated scenario instead of radiation transfer, i.e. the photon diffusion with constant opacity. In general, the evolution of the ejecta’s internal energy could be written as
@@
For the convenience of numerical calculation, we discretize the ejecta into $N$ shells and number the shells from $n=0$ to $n=N$. An arbitrary physical quantity $Q$ for a certain shell $n$ is denoted as $\delta Q_{n}$. Sometimes we drop the subscript $n$ when referring to any shell in general. For each shell, the evolution of internal energy follows
@@
where $\delta L_{\rm out}^{\prime}$ represents the net outflow of the internal energy into this shell due to the photons’ diffusion.
Suppose every shell of the merger ejecta initially has a specific bulk motion velocity $v$, in the lab frame. The velocity difference between two adjacent shells is denoted as $\delta v$. Assume the density of the ejecta depends on the velocity, which is given by
@@
Assuming the ejecta is isotropic, the volume of a shell can be calculated as
@@
The radioactive power released through nuclear reaction has been detailed studied in the literature (e.g. Metzger et al., 2010; Korobkin et al., 2012; Rosswog et al., 2018; Zhu et al., 2021; Wu et al., 2022; Chen & Liang, 2024). Here we adopt the analytical proposed by Korobkin et al. (2012). For unit mass, the power can be described as a function of time, which reads as
@@
$\epsilon_{\rm th}(t,x)=0.36\left[{\rm exp}(-aX)+\frac{{\rm ln}(1+2bX^{d})}{2bX^{d}}\right]$ (12)
@@
The interaction between the post-merger NS wind and the merger ejecta can excite a forward shock (FS) in the ejcta, which can both alter the dynamics of the ejecta and provide an additional heat source. The shock heating rate is denoted as $L_{\rm sh}^{\prime}$ and should be included into the extra heating term $L_{\rm ext}^{\prime}$. The physical scenario and basic formulae for the dynamics of engine-fed kilonova have been discussed in detail in Paper I. In this work, the setups are generally identical to those in Paper I, with only a few slight differences.
@@
In Paper I, we assumed that the bulk motion velocity and the magnetization parameter for the unshocked NS wind are constant in both time and space, as they are believed not to significantly influence the energy injection process, especially when they were sufficiently large. The luminosity of the unshocked NS wind was also assumed to be constant as our primary focus was on examining the dependence of energy injection efficiency on the wind’s luminosity. However, in this paper, while maintaining the same setups for the bulk motion and magnetization of the unshocked wind, we introduce a more realistic approach where the luminosity evolves in accordance with the spindown of the central NS, which is given by
$L_{w}(t)=\eta L_{\rm sd,0}\left(1+\frac{t}{t_{\rm sd}}\right)^{-2}$ (13)
where $t_{\rm sd}=E_{\rm tot}/L_{\rm sd,0}$ is the spin-down timescale of the NS due to magnetic dipole radiation, with $E_{\rm tot}$ as the total energy budget of the NS wind and $L_{\rm sd,0}$ as the initial spin-down luminosity. A beaming factor $\eta$ is introduced, as the fast rotation of the central NS can focus roughly $90\%$ of the spin-down power into the direction of the spin axis (Wang et al., 2024c). Therefore, $\eta=0.1$ is adopted here in the scenario with nearly isotropic merger ejecta. Regarding the ejecta, we introduced a gradient for the initial bulk-motion velocity, as discussed in the beginning of Section 2. Instead of manually setting up the initial position ($r_{0}$) of the contact discontinuity between the wind and the ejecta, as we did in Paper I, we assume a time delay between the ejecta launching and the generation of the NS wind ($\delta t_{l}$; in the central engine frame)  [^1] The results are not sensitivity to the exact value of $\delta t_{l}$, as long as it is much shorter than the spin-down timescale of the post-merger magnetar. Then, we can calculate the time of the first contact between the NS wind and the ejecta, which reads as
@@
When $\beta_{\rm min}=0.05$, $t_{c}=1.05(\delta t_{l}/1{\rm s})~{}{\rm s}$ and $r_{0}=1.57\times 10^{9}(\delta t_{l}/1{\rm s})~{}{\rm cm}$.
For the unshocked ejecta, we assume that the bulk-motion velocity of each shell remains constant over time. The velocity of the material immediately preceding the FS can be simply calculated as $v_{f}=r_{f}/t$, where $r_{f}$ is the radius of the FS and $t$ is time in the lab frame. The density upstream of the FS should be  [^2] To maintain consistency with Paper I, the symbol “$\rho$” will henceforth represent the density in the rest frame of the local fluid, without the need for a superscript $\prime$”.
@@
$+\frac{dP_{\rm sph}^{\prime}}{dr_{d}}\bigg{|}_{\rm ra}+\frac{dP_{\rm sph}^{\prime}}{dr_{d}}\bigg{|}_{e}+\frac{dP_{\rm sph}^{\prime}}{dr_{d}}\bigg{|}_{\rm diff}$
The first three terms on the right-hand side of Equation 18 represent the accumulation of $P_{\rm sph}$ due to the reverse and forward shocks, as well as the effect of adiabatic expansion, respectively. These topics have been discussed in detail in Paper I. Here, we clarify the physical meaning of the last three terms and derive their expressions.
@@
Similarly, the term related to emission is expressed as
@@
The luminosities for photon diffusion across different layers ($L_{\rm diff}^{\prime}$) and of thermal emission ($\delta L_{e}^{\prime}$) are detailed studied in Section 2.4 and Section 2.5, respectively.
@@
The pressure in each shell is thus $p=(1/3)(\delta E_{\rm int}^{\prime}/\delta V^{\prime})$. Considering a long-lived NS (magnetar) as the central remnant of a BNS merger, the physical quantities preceding the RS, i.e. the wind density and magnetic pressure, can be derived directly from the wind luminosity at a given time and the position of the reverse shock ($r_{r}$). Assuming a constant Lorentz factor for the bulk motion of the NS wind ($\Gamma_{w}$), these derivations have been elaborated in Paper I. The results remain insensitive to the precise value of $\Gamma_{w}$ when it is sufficiently large, and thus $\Gamma_{w}=1000$ is adopted in this study.
@@ 2.3 X-ray irradiation
In Paper I, we assumed that the RS crosses the mangetar’s wind and vanishes within a very short timescale due to the wind’s high magnetization level, preventing the formation of a long-lasting pulsar wind nebula. However, some authors argued that the magnetar’s wind is not purely magnetic flux, but also contains a large fraction of electron-positron pairs. The synchrotron radiation emitted by the RS-accelerated pairs could convert the wind’s magnetic energy into X-ray photons (Metzger & Piro, 2014; Wang et al., 2024a). Furthermore, the self-dissipation of the magnetic field may be another channel for producing X-ray photons (Zhang & Yan, 2011). Regardless, these X-ray photons would be absorbed by the merger ejecta and contribute to the kilonova emission. In this study, although the NS wind is assumed to be dominated by the magnetic field, we do not completely ignore the potential X-ray irradiation. We introduce another efficiency parameter, $\eta_{X}$, to represent the fraction of wind luminosity that can be converted into X-ray radiation. The X-ray heating rate, denoted as $L_{X}^{\prime}$, should also be included in the extra heating term $L_{\rm ext}^{\prime}$.
@@
where $L_{X,t}^{\prime}=\xi_{X}\eta_{X}L_{w}/\Gamma^{2}$. The subscript ’$1$’ and ’$f$’ in $k_{1,f}$ indicate that the X-ray photons propagate from the innermost shell (the $1$st shell) in the ’forward’ direction. $k_{1,f}$ is obtained by modeling the photon diffusion, by equaling the timescale for the photons diffusing from the innermost layer to the $k_{1,f}$th layer, with the time resolution for our calculation. This will be detailed presented in Section 2.4. Similar as Equation 24, a thermalization factor $\xi_{X}$ is introduced. Again, it is expressed as
@@
where $\tau_{\rm ej}$ is the optical depth for photons diffusing from the innermost to the outermost layers. For $\eta_{X}\ll 1$, the dissipation of magnetic field would not influence the dynamics of the system, so that we will not modify the formulae related to the mechanical model of the blastwave. When $\eta_{X}\sim 10^{-2}$, the contribution of X-ray irradiation would be comparable with the shock heating, because the shock heating efficiency obtained from Paper I is also $\xi_{t}\sim 0.01$.

### v2

@@ 1 Introduction
Kilonova has garnered significant attention in recent years, particularly since the detection of AT2017gfo in association with the binary-neutron-star (BNS) merger GW170817 (e.g. Arcavi et al., 2017; Chornock et al., 2017; Cowperthwaite et al., 2017; Drout et al., 2017; Evans et al., 2017; Gao et al., 2017a; Kasen et al., 2017; Kilpatrick et al., 2017; Nicholl et al., 2017; Shappee et al., 2017; Smartt et al., 2017; Tanvir et al., 2017; Villar et al., 2017; Metzger, 2019, and the references there in, for a review). Other kilonova candidates have been observed in association with short-duration gamma-ray bursts (GRBs), which are believed to be also of a neutron star merger origin (e.g. Tanvir et al., 2013; Berger et al., 2013; Fan et al., 2013; Yang et al., 2015; Jin et al., 2016; Gao et al., 2015, 2017b). Most recently, some unusual GRB events, specifically GRB 211211A and GRB 230307A, have long duration but with kilonova-like emission signal associated (Rastinejad et al., 2022; Yang et al., 2022; Troja et al., 2022; Gao et al., 2022; Zhu et al., 2022; Yang et al., 2023; Gillanders et al., 2023; Levan et al., 2023). The James Webb Space Telescope’s mid-infrared follow-up of GRB 230307A revealed an emission line at 2.15 microns in the spectrum, which is interpreted as tellurium (Levan et al., 2023; Gillanders et al., 2023). Additionally, the rapidly decaying bolometric luminosity of the kilonova at the later stage supports the presence of lanthanide in the ejecta. This kilonova signal is hence generally believed to be powered by r-process radioactive heating (Yang et al., 2023).
Notably, the presence of a plateau in the X-ray afterglow light curve of GRB 230307A, detected by the Lobster Eye Imager for Astronomy (LEIA), a pathfinder of the Einstein Probe mission, indicates that the merger remnant is a magnetar (Sun et al., 2023; Ling et al., 2023). Previous studies have speculated that a post-merger magnetar could drive an engine-fed kilonova (mergernova), which is anticipated to be significantly brighter than a typical r-process kilonova (Yu et al., 2013; Metzger & Piro, 2014; Gao et al., 2015; Ren et al., 2019; Sarin et al., 2022; Wang et al., 2024a). Therefore, the detection of the engine-fed kilonova signal serves as another crucial piece of evidence regarding the production of a long-lived magnetar. In reality, by adjusting the free parameters in the theoretical model of an engine-fed kilonova, the predicted luminosity of thermal radiation from the merger ejecta can span several orders of magnitude. When assuming a relatively low spin-down luminosity and low heating efficiency, the peak luminosity of the engine-fed kilonova can be reduced to approximately $10^{42}~{}{\rm erg/s}$ (e.g. Li et al., 2018; Yu et al., 2018; Ren et al., 2019; Yang et al., 2022, for the kilonovae associated with GW170817 and GRB 211211A), which can also be achieved within the framework of a typical r-process kilonova, given a reasonably large ejecta mass. Hence, for certain critical scenarios, it becomes extremely challenging to discern scenarios with or without energy injection. In some previous studies, a pulsar wind nebula (PWN) has been introduced by considering the reverse shock (RS) propagating in a magnetar wind dominated by electron-positron pairs (Metzger & Piro, 2014; Ren et al., 2019; Murase et al., 2018; Wang et al., 2024a). They have comprehensively modeled the generation of non-thermal radiations from the PWN and how these non-thermal radiations are absorbed by the ejecta. In their models, the PWN does not significantly alter the dynamics of the merger ejecta, thus the peak time of the kilonova might remain roughly unchanged. A considerable portion of the non-thermal photons would escape from the ejecta, serving as compelling evidence for the presence of a post-merger magnetar (Murase et al., 2018).
@@
We have made continuous efforts to more precisely model the engine-fed kilonova. Our first paper in this series (Ai et al., 2022) provided a detailed study for the dynamics of the ejecta resulting from the interaction between the magnetar wind and the merger ejecta. It is under the framework of the mechanical model for magnetized relativistic blastwaves (Ai & Zhang, 2021a). For convenience, we henceforth refer to Ai et al. (2022) as Paper I. In Paper I, we found that a significant fraction ($\xi$ as high as $\sim 0.6$) of the energy from the post-merger magnetar’s wind can be converted into the kinetic energy of the ejecta’s bulk motion. Meanwhile, the forward shock excited in the ejecta provides an additional heating source during its propagation, with a heating efficiency of $\xi_{t}\sim(0.006-0.3)$. In this work, building upon the dynamical model proposed in Paper I, we focus on radiation. The basic formula for modeling both the heating and radiation processes of the engine-fed kilonova is presented in Section 2. The results, including the multi-band light curves and the spectral energy distribution (SED), are shown in Section 3. The conclusion and discussion are detailed in Section 4.
@@ 2 Methods
The interaction of magnetar wind and the merger ejecta excites a forward shock (FS) in the ejecta and a reverse shock (RS) in the highly magnetized wind. A blastwave typically refers to the region situated between the FS and RS. The magnetization parameter ($\sigma_{w}$) of the wind likely remains significantly greater than unity all the way into the magnetar magnetosphere. The RS would propagate rapidly through the wind and quickly vanish near the magnetosphere, where the pressures of the unshocked wind and the blastwave become comparable (Zhang & Kobayashi, 2005; Ma & Zhang, 2022). We set this location as the inner boundary of for our dynamics calculation and consider the region extending from this inner boundary to the FS as the blastwave. In Paper I, we studied in detail the dynamics of this system and the heating of the ejecta due to the FS. In that paper, the unshocked ejecta was assumed to be wind-like, with a density profile of $\rho_{\rm ej}\propto r^{-2}$, where $r$ represents the distance from the central engine in the lab frame. We also assumed that the ejecta is expanding like a shell, where the speed along radial direction is constant throughout the space. Additionally, since radioactive heating and radiation do not significantly influence the dynamics, we neglected both of them, as well as photon diffusion inside the ejecta. Therefore, the evolution of the internal energy of the blastwave (equivalent to the evolution of integrated gas pressure $P_{\rm sph}$ as described in Paper I) depends solely on shock heating and the adiabatic expansion of the ejecta.
In this paper, to model the radiation of an engine-fed kilonova, we must comprehensively consider the evolution of the ejecta’s internal energy. Generally, this evolution can be expressed as (Yu et al., 2013). [^1] For intuitive purposes, we directly present the evolution of internal energy here, rather than the integrated gas pressure $P_{\rm sph}$. The equation describing the evolution of $P_{\rm sph}$ should be modified accordingly, as demonstrated in Section 2.2, where the mechanical model for blastwaves is introduced.
@@
For the convenience of numerical calculation, we discretize the ejecta into $N$ shells and number the shells from $n=1$ to $n=N$. An arbitrary physical quantity $Q$ for a certain shell $n$ is denoted as $\delta Q_{n}$. Sometimes we drop the subscript $n$ when referring to any shell in general. For each shell, the evolution of internal energy follows
@@
where $\delta L_{\rm out}^{\prime}$ represents the net outflow of the internal energy from a certain shell due to photon diffusion. The details for photon diffusion across different shells will be discussed in Section 2.4.
Another difference with Paper I is that here we introduce a gradient for bulk-motion velocity of the unshocked ejecta, to more realistically describe the outflow properties. Suppose every shell of the merger ejecta initially has a specific bulk motion velocity $v$, in the lab frame. The velocity difference between two adjacent shells is denoted as $\delta v$. We assume the density of the ejecta depends on the velocity, which is given by
@@
Also assuming the ejecta is isotropic, the volume of a shell can be calculated as
@@ 2.1 Radioactive heating
The radioactive power released through nuclear reaction has been studied in detail in the literature (e.g. Li & Paczyński, 1998; Metzger et al., 2010; Korobkin et al., 2012; Rosswog et al., 2018; Zhu et al., 2021; Wu et al., 2022; Chen & Liang, 2024). Here we adopt the analytical treatment proposed by Korobkin et al. (2012). For unit mass, the power can be described as a function of time, which reads
@@
$\epsilon_{\rm th}(t,x)=0.36\left[{\rm exp}(-aX)+\frac{{\rm ln}(1+2bX^{d})}{2bX^{d}}\right],$ (12)
@@ 2.2 Shock heating
The interaction between the post-merger NS wind and the merger ejecta can excite an FS in the ejcta, which can both alter the dynamics of the ejecta and provide an additional heating source. The shock heating rate is denoted as $L_{\rm sh}^{\prime}$ and should be included into the extra heating term $L_{\rm ext}^{\prime}$. The key to studying shock heating lies in solving the dynamics of the blastwave, as detailed in Paper I.
@@
In Paper I, we assumed that the bulk motion velocity and the magnetization parameter for the unshocked NS (magnetar) wind are constant in both time and space, as they are believed not to significantly influence the energy injection process, especially when they were sufficiently large. The luminosity of the unshocked NS wind was also assumed to be constant as our primary focus was on examining the dependence of energy injection efficiency on the wind’s luminosity. However, in this paper, while maintaining the same setups for the bulk motion and magnetization of the unshocked wind, we introduce a more realistic approach where the luminosity evolves in accordance with the spindown of the central NS, which is given by
$L_{w}(t)=\eta L_{\rm sd,0}\left(1+\frac{t}{t_{\rm sd}}\right)^{-2},$ (13)
where $t_{\rm sd}=E_{\rm tot}/L_{\rm sd,0}$ is the spin-down timescale of the NS due to magnetic dipole radiation, with $E_{\rm tot}$ being the total energy budget of the NS wind and $L_{\rm sd,0}$ being the initial spin-down luminosity. We assume that the wind is generated and accelerated near the light cylinder of the newly-born millisecond magnetar at $r_{\rm acc}=10^{7}{\rm cm}$. Beyond this radius, both the Lorentz factor for the bulk motion of the NS wind ($\Gamma_{w}$) and its magnetization parameter ($\sigma_{w}$) are considered constant in both space and time for the unshocked wind. As has been elaborated in Paper I, the results remain insensitive to the exact value of $\Gamma_{w}$ and $\sigma_{w}$ when they are sufficiently large, and thus $\Gamma_{w}=10^{3}$ and $\sigma_{w}=10$ are adopted in this study. A beaming factor $\eta$ is introduced, as the fast rotation of the central NS can focus roughly $90\%$ of the spin-down power into the direction of the spin axis (Wang et al., 2024c). Therefore, $\eta=0.1$ is adopted here in the scenario with nearly isotropic merger ejecta. The physical quantities preceding the RS, i.e. the wind density and magnetic strength, can be derived directly from the wind luminosity at a given time and the position of the RS ($r_{r}$) (See Equations 31 and 32 in our Paper I). The RS in the wind reaches the magnetosphere of the magnetar quickly after it was exicited and subsequently dissipates there (set as inner boundary). Then $\Gamma_{w,{\rm acc}}=\Gamma$ should be satisfied, where $\Gamma_{w,{\rm acc}}$ and $\Gamma$ represent the Lorentz factor for the bulk motion of the unshocked wind at $r_{\rm acc}$ and the bulk motion of the blastwave, respectively. With $r_{r}=r_{\rm acc}$ and $\Gamma_{w}=\Gamma$, and assuming a constant $\sigma_{w}$, the physical quantities of the wind at the inner boundary, where the wind is generated, can be calculated at each time.
To study the blastwave system, one should, in principle, set the initial point of contact between the magnetar wind and the merger ejecta. In Paper I, we designated the initial position of the contact discontinuity between the wind and ejecta as $r_{0}$. Physically speaking, this is a catch-up problem between the wind and the ejecta, as the ejecta is launched during the merger while the continuous wind is generated after the central rigidly rotating NS has stabilized. It is believed that a post-merger NS will experience the differentially rotating phase within at least several hundred millisecond after the merger (e.g. Radice et al., 2018). Therefore, here we assume a time delay between the ejecta launching and the generation of the constant NS wind ($\delta t_{l}$; in the central engine frame). Then, we can calculate the time of the first contact between the NS wind and the ejecta, which reads as
@@
When $\beta_{\rm min}=0.05$, $t_{c}=1.05(\delta t_{l}/1{\rm s})~{}{\rm s}$ and $r_{0}=1.57\times 10^{9}(\delta t_{l}/1{\rm s})~{}{\rm cm}$. Although the exact value of $t_{l}$ is unknown, since the magnetar wind lasts for a duration much longer than $t_{l}$, the value of $t_{l}$ is unlikely to have a substantial influence on the results. We tested with $t_{l}=0.1$s, $1$s, and $10$s, respectively, and found little difference in the outcomes. Consequently, we use $t_{l}=1$s consistently throughout the paper for simplicity.
For the unshocked ejecta, we introduced a gradient for the initial bulk-motion velocity, as discussed in the beginning of Section 2, and assume that the bulk-motion velocity of each shell remains constant over time. The velocity of the material immediately preceding the FS can be simply calculated as $v_{f}=r_{f}/t$, where $r_{f}$ is the radius of the FS and $t$ is time in the lab frame. The density upstream of the FS should be [^2] To maintain consistency with Paper I, the symbol “$\rho$” will henceforth represent the density in the rest frame of the local fluid, without the need for a superscript $\prime$”.
@@
$+\frac{dP_{\rm sph}^{\prime}}{dr_{d}}\bigg{|}_{\rm ra}+\frac{dP_{\rm sph}^{\prime}}{dr_{d}}\bigg{|}_{e}+\frac{dP_{\rm sph}^{\prime}}{dr_{d}}\bigg{|}_{\rm diff},$
where $dr_{d}=\beta dt$ represents the evolution of the position of the contact discontinuity, and $\beta$ is the dimensionless velocity of the blastwave’s bulk motion. The first three terms on the right-hand side of Equation 18 represent the accumulation of $P_{\rm sph}$ due to the reverse and forward shocks, as well as the effect of adiabatic expansion, respectively. These topics have been discussed in detail in Paper I. Here, we clarify the physical meaning of the last three terms and derive their expressions.
@@
Here $v_{f}$ ($\beta_{f}$) represents the dimensionless speed for the propagation of the forward shock, and $\hat{\gamma}=4/3$ is adopted as the adiabatic index for the matter in the blastwave. Similarly, the term related to radiation is expressed as
@@
The luminosities for photon diffusion across different layers ($L_{\rm diff}^{\prime}$) and of thermal radiation ($\delta L_{e}^{\prime}$) are detailed studied in Section 2.4 and Section 2.5, respectively.
@@
The pressure in each shell is thus $p=(1/3)(\delta E_{\rm int}^{\prime}/\delta V^{\prime})$.
@@ 2.3 X-ray irradiation
In Paper I, we assumed that the RS crosses the mangetar’s wind and vanishes within a very short timescale due to the wind’s high magnetization, preventing the formation of a long-lasting pulsar wind nebula. However, some authors argued that the magnetar’s wind is not purely magnetic flux, but also contains a large fraction of electron-positron pairs. The synchrotron radiation emitted by the RS-accelerated pairs could convert the wind’s magnetic energy into X-ray photons (Metzger & Piro, 2014; Wang et al., 2024a). Furthermore, the self-dissipation of the magnetic field may be another channel for producing X-ray photons (Zhang & Yan, 2011). Regardless, these X-ray photons would be absorbed by the merger ejecta and contribute to the kilonova radiation. In this study, although the NS wind is assumed to be dominated by the magnetic field, we do not completely ignore the potential X-ray irradiation. We introduce another efficiency parameter, $\eta_{X}$, to represent the fraction of wind luminosity that can be converted into X-ray radiation. The X-ray heating rate, denoted as $L_{X}^{\prime}$, should also be included in the extra heating term $L_{\rm ext}^{\prime}$.
@@
where $L_{X,t}^{\prime}=\xi_{X}\eta_{X}L_{w}/\Gamma^{2}$. The subscript “$1$”and “$f$” in $k_{1,f}$ indicate that the X-ray photons propagate from the innermost shell (the $1$st shell) in the ’forward’ direction. $k_{1,f}$ is obtained by modeling the photon diffusion, by equaling the timescale for the photons diffusing from the innermost layer to the $k_{1,f}$th layer, with the time resolution for our calculation. This will be detailed presented in Section 2.4. Similar as Equation 24, a thermalization factor $\xi_{X}$ is introduced. Again, it is expressed as
@@
where $\tau_{\rm ej}$ is the optical depth for photons diffusing from the innermost to the outermost layers.